A disk between vertebrae in the spine is subjected to a shearing force of 825 N. Find its shear deformation, taking it to have a shear modulus of 1.20 Ý 109 N / m2. The disk is equivalent to a solid cylinder 0.900 cm high and 3.00 cm in diameter.
Answer #1

youngs modulus=(F*L)/(A*delL)

shear deformation1.2*109=825*0.009/pir2

shear deformation=8.753*10-6
Answer #2

shear deformation ==

shear deformation = (1/(1.2*10^9))*(825/(pi*(0.015)^2)) *0.09 = 0.00008753521 m